In a triangle ABC, ( b+c+-a) tan A\2 is equal to a) 2triangle\s. B) triangle \s C) triangle s\bc. D) s/aR

In a triangle ABC, (b+c+-a)  tan A\2 is equal  to
a) 2triangle\s.   B) triangle \s
C) triangle  s\bc.    D) s/aR

Grade:11

1 Answers

Vikas TU
4 years ago
Hiii 
I a triangle ABC (b+c-a) tan A/2
First  write Tan (A/2) , Tan (B/2) , Tan (C/2) formula in terms of s
Then apply heroins formulae, 
at last put the value of S = (a+b+c)/2
You will get 
(b+c-a)Tan (A/2) = (a+c-b) Tan(B/2) = (a+c-b) Tan(C/2)